# Crepe: A Mobile Screen Data Collector Using Graph Query

### v1

@@ E $\rightarrow$ S;
Specifically, Graph Queries encapsulate combinations of the screen entity relations for the target UI element and Crepe creates combinations can uniquely identify only the target UI element on screen (Figure 3 (2)). The combinations of screen entities are empirically set through experimentation. All graph queries will have hasPackageName and hasClassName to establish the context. In most cases, more than one graph query can uniquely locate the target UI element on the screen. In these scenarios, to ensure the final selected Graph Query reflects the user’s intention of collecting the target data, we used a large language model [^4] We used GPT-3.5-turbo in our implementation due to its ease of access, short inference time, and decent performance observed on our task., to select the Graph Query that best matches users’ description of target data. The description of the target data to collect was an input field in Crepe’s graphical user interface (the details of our prompt can be found in Appendix A.1). Altogether, Crepe can effectively generate graph queries that can reliably, uniquely identify and locate the target UI element on screen.
@@ 3.2.3. Graph Query execution on UI Snapshots: retrieving the target data
After a graph query is created, it can be used to collect target data on any UI Snapshot with elements sharing the same set of entity relations. In the background, Crepe runs the Graph Query every time a relevant UI snapshot is updated by changes in screen content (Figure 3 Section (B)). We also limit the Graph Query to only run in the target app package, in order to effectively reduce unnecessary battery usage of Crepe. After idenfitying and locating the target UI element on screen (3), Crepe can collect the target data to collect through the element’s attributes such as text and content description.
@@ 3.2.4. Summary
Graph Query showcases three major strengths in screen data collection. First, it provides much more flexibility in identifying and locating target UI elements through their implementation attributes, semantic content, hierarchical screen structure position, and screen location relations (3.2.1). For instance, a researcher can utilize the pre-defined hasPrice relation to collect Uber drivers’ task payment information, a common task on gig worker support platforms (Calacci and Pentland, 2022). In the future, developers can also extend existing entity relation types for more customized data collection tasks. Second, the data collection mechanism using Graph Query is fully deterministic and more reliable than alternative solutions like using Optical Character Recognition (OCR) or deep-learning based detection models. Lastly, by utilizing Android’s Accessibility Service, using graph query can easily integrate the collection of user interaction data, a feature included in Crepe yet unsupported in most vision-based data collection tools. In the future, Crepe can be extended to collect more rich media information such as screenshots under the study participants’ consent555 Link to the documentation of Android Accessibility Service’s API for taking screenshots..
@@ 3.3. Implementation
Crepe was implemented using Android, Google Firebase Realtime Database, OAuth, and OpenAI API. The app was developed in Java using Android Studio and is compatible with mobile devices running Android 9.0 or above, which in March 2024 takes up around 90% of the Android mobile phone market. Firebase Realtime Database was selected due to its high-standard data encryption both in transit and at rest, real-time data synchronization, and reliable user authentication. To implement the query language, Crepe leverages the Accessibility Service and captures Accessibility Events as a result of changes in the content of the screen. The captured Accessibility Event contains the new UI screen hierarchy [^6] https://developer.android.com/reference/android/view/accessibility/AccessibilityEvent. The code for Crepe will be open-sourced to enable easy adoption and further development within the research community.
4. Example Use Case and User Testing
Participant ID Mobile OS Device Period Battery Usage
P1 Android 14 Pixel 8 (own) 72 hours 1%
P2 Android 14 Pixel 6 (provided) 72 hours 14%
P3 Android 14 Pixel 6 (provided) 72 hours 9%
P4 Android 13, OxygenOS 13.1 OnePlus 9R (own) 24 hours 0.91%
P5 Android 12, Oxygen OS 12.1 OnePlus Nord (own) 24 hours N/A
P6 Android 14 Pixel 6 (provided) 24 hours 2%
P7 Android 9 Galaxy S10 (own) 24 hours 0%
Table 2. Mobile device usage data for study participants over 24 to 72 hour periods, including mobile OS, device model, and battery usage percentage. Three of the participants did not have an Android phone and used our provided device, with which they were asked to use Instagram following their regular habits during the study period. The participants decided the time period of the study based on their time schedule. Note that P5 reported that they could not found the battery usage of Crepe after the collection. We were not able to help them find it either, despite our best efforts.
In this section, we describe a potential use case of collecting Instagram Story advertisements to demonstrate the collection mechanism and utility of Crepe. We also deployed a user study to understand Crepe’s collection performances and limitations in the wild.
4.1. Example Use Case: Collecting Instagram Story Advertisements
Instagram Story advertisements are sponsored short image/video segments that appear between an Instagram user’s following content. They usually come with a “Sponsored” label to be distinguished from users’ regular following content. Mutiple research studies have been conducted in recent years around Instagram Story advertisements, to understand how users perceive and understand this specific form of advertising (Balaban et al., 2022; Haldborg Jørgensen et al., 2023; Fakhradyan, 2021). However, our the best of our knowledge, no academic research to date has been able to collect a dataset of Instagram story advertisements. Such data is also not available from Instagram’s developer API  [^7] https://developers.facebook.com/docs/instagram-api/.
We believe that getting access to a dataset of Instagram’s Story advertisements for individual users can enable research contributions in fields including social media research (Hanbazazh and Reeve, 2021), algorithm auditing (Ulbricht and Yeung, 2022; Eslami et al., 2015, 2016), and usable privacy (Chen et al., 2024). Specifically, we can empirically observe users’ interactions with Instagram Story ads, as well as inferring dwelling time over Story ads based on the first and last content change while a specific ad appears on screen. More generally, some broader research questions we can answer with the dataset include:
@@ 4.1. Example Use Case: Collecting Instagram Story Advertisements
    What empirical understanding of Instagram’s personalized advertisement algorithms can we obtain from quantitative data analysis?
@@
    How Instagram Story advertisements reveal the platform’s portraits of individual users’ preferences?
@@
    What are users’ folk theories of Instagram’s personalized advertisement algorithm?
4.2. User Testing
We conducted a user study to evaluate the performance and limitations of Crepe in real-world scenarios. It is important to note that the primary objective of this study was not to collect high-quality, immediately usable datasets to support future research, but rather to assess the real-world performance of our data collector and its impact on participants’ regular mobile device usage experience. To this end, the study aimed to address the following research questions:
  1. (1)
@@ 4.2. User Testing
    RQ1: How does Crepe perform in a real-world data collection scenario?
  2. (2)
    RQ2: What limitations does Crepe have when collecting data in the wild?
  3. (3)
    RQ3: How does Crepe impact the regular everyday mobile phone device usage experience of participants?

### v2

@@ E $\rightarrow$ S;
[…]
This figure presents a table outlining the context-free grammar (CFG) rules used to construct Graph Queries in the Crepe system. The table has two columns: the left column lists grammar expressions, and the right column shows their corresponding production rules.
  • E can be replaced by either ‎‘e;‘ or ‎‘S;‘
  • S can be replaced by ‎‘(join r E)‘ or ‎‘(and S S)‘
  • T can be replaced by ‎‘(ARG_MAX r S)‘ or ‎‘(ARG_MIN r S)‘
  • Q, the initial non-terminal symbol, can be replaced by either ‎‘S‘ or ‎‘T‘
Specifically, Graph Queries encapsulate combinations of the screen entity relations for the target UI element and Crepe creates combinations can uniquely identify only the target UI element on screen (Figure 4, Step 2). The combinations of screen entities are empirically set through experimentation. All Graph Queries will have hasPackageName and hasClassName to establish the context. In most cases, more than one Graph Query can uniquely locate the target UI element on the screen.
3.2.3. User Intent Disambiguation: Selecting the Best Graph Query
As described above, for each target data, Crepe generates a few Graph Query candidates that can each uniquely match the target data. However, some candidates are more generalizable to other screens—for example, to collect Instagram story ads, a query that collects the text above “Sponsored” is more likely to collect all target data, than a query that collects text that shows up at a specific screen location (the first and third queries in Figure 4, Step 2). We created a set of heuristics to rank the generalizability of generated queries; however, through our user studies, we learned that it is best to engage the researcher in this selection process.
Through iterative design with researchers (details in Section 4.2) and taking inspirations from past research (Li et al., 2018), our final design presents the top Graph Query candidates in natural language phrases for the user to select from. We translate Graph Queries into natural language phrases, such as “the UI element above the Button that says Sponsored”, to show to the user instead of its original query language form (Figure 4, Section 2, option 1). We translated the Graph Queries so that the researchers, especially the non-technical ones, are not exposed to unnecessary technical details of Crepe. We translate the Graph Queries using a large language model [^4] We used GPT-3.5-turbo in our implementation due to its ease of access, short inference time, and empirically high-quality outputs observed on our task. Our prompt can be found in Appendix A.1.. This ensures the selected Graph Query best reflects the researcher’s data collection goals.
@@ 3.2.4. Graph Query execution on UI Snapshots: retrieving the target data
After a Graph Query is created, it can be used to collect target data on any UI Snapshot with elements sharing the same set of entity relations. In the background, Crepe runs the Graph Query every time a relevant UI snapshot is updated by changes in screen content (Figure 4, Section B). We also limit the Graph Query to only run in the target app package, in order to effectively reduce unnecessary battery usage of Crepe. After identifying and locating the target UI element on screen (Step 3), Crepe can collect the target data to collect through the element’s attributes such as text and content description.
@@ 3.2.5. Summary
Graph Query showcases three major strengths in screen data collection. First, the data collection mechanism using Graph Query is fully deterministic and thus much more reliable than alternative solutions like using Optical Character Recognition (OCR) or deep-learning based detection models. Second, Graph Query provides much more flexibility in identifying and locating target UI elements through their implementation attributes, semantic content, hierarchical screen structure position, and screen location relations (Section 3.2.1). For instance, a researcher can utilize the pre-defined hasPrice relation to collect Uber drivers’ task payment information, a common task on gig worker support platforms (Calacci and Pentland, 2022). In the future, developers can also extend existing entity relation types for more customized data collection tasks. Lastly, by utilizing Android’s Accessibility Service, using Graph Query can easily integrate the collection of user interaction data, a feature included in Crepe yet unsupported in most vision-based data collection tools. In the future, Crepe can be extended to collect more rich media information such as screenshots under the study participants’ consent555 Link to the documentation of Android Accessibility Service’s API for taking screenshots..
@@ 3.3. Implementation
Crepe was implemented using Android, Google Firebase Realtime Database, OAuth, and OpenAI API. Note that we only used OpenAI large language model for Graph Query translation (Section 3.2.3) and the whole data collection process does not involve any deep learning. The app was developed in Java using Android Studio and is compatible with mobile devices running Android 9.0 or above (around 95.4% of the Android mobile phone market as of February 2025666https://composables.com/android-distribution-chart). Firebase Realtime Database was selected due to its high-standard data encryption both in transit and at rest, real-time data synchronization, and reliable user authentication. To implement the query language, Crepe leverages the Accessibility Service and captures Accessibility Events as a result of changes in the content of the screen. The captured Accessibility Event contains the new UI screen hierarchy [^7] https://developer.android.com/reference/android/view/accessibility/AccessibilityEvent. The code for Crepe will be open-sourced to enable easy adoption and further development within the research community.
4. Evaluations
To test and validate Crepe, we conducted a series of three evaluation studies. The first two studies focused on researchers’ experience to collect data (Study 1) and Crepe’s data collection accuracy (Study 2). The third study investigated Crepe’s battery usage and perceived interruptions in the wild for participants (i.e. data contributors) (Study 3). The first two studies were designed to be in-lab, while Study 3 was conducted as a field study. These three studies together aim to answer the following research questions:
@@ 4. Evaluations
    RQ1: How well does Crepe support researchers in collecting mobile screen data (Study 1)?
@@
    RQ2: How accurately does Crepe collect target data on participants’ screens (Study 2)?
@@
    RQ3: How does Crepe influence device performance for participants (i.e. data contributors) in a real-world data collection scenario (Study 3)?
4.1. Evaluation Scenarios
We wanted our performance evaluation of Crepe to cover a range of different data collection needs. Looking broadly into previous research in HCI, CSCW, and Data Science, we extracted three scenarios where researchers can use Crepe to collect user screen data, Our evaluation studies were based in these scenarios.
Figure 5. We identified three potential usage scenarios of Crepe for our evaluation. Scenario 1 collects the advertiser information in Instagram Story ads, scenario 2 collects Uber ride pricing when user requests a ride, and scenario 3 collects the top feed item in Chrome Discover Feed recommendations. Crepe shows a yellow highlight overlay every time the target data is collected, providing full transparency to the collection participants.
[…]
This figure shows three mobile phone screenshots arranged horizontally, each labeled as a different scenario. In Scenario 1 (left), an Instagram Story advertisement is shown with the text “Inspiration is everywhere. Find yours.” along with several colorful images featuring someone in casual settings. In Scenario 2 (middle), an Uber app interface displays ride options near a Walmart Supercenter. A yellow highlight box surrounds ”UberXL” pricing information showing ”7 min away” and ”Affordable rides for groups up to 6.” In Scenario 3 (right), a Google Chrome mobile browser interface shows the Discover feed. A yellow highlight box surrounds a news article preview titled “This Is A Big Deal - Bitcoin And Crypto Now Braced For A Huge U.S. Price Earthquake” from Forbes published 7 hours ago. Each screenshot represents a different type of digital data that could be collected for research purposes, with the yellow highlighting indicating the specific elements of interest.
Scenario 1: Instagram Story Advertisements
Instagram Story advertisements are sponsored short image/video segments that appear between an Instagram user’s following content. They usually come with a “Sponsored” label to be distinguished from users’ regular following content. Mutiple research studies have been conducted in recent years around Instagram Story advertisements, to understand how users perceive and understand this specific form of advertising (Balaban et al., 2022; Haldborg Jørgensen et al., 2023; Fakhradyan, 2021). However, it is not easy to collect a dataset of Instagram story advertisements.
Getting access to a dataset of Instagram’s Story ads can enable research contributions in fields including social media research (Hanbazazh and Reeve, 2021), algorithm auditing (Ulbricht and Yeung, 2022; Eslami et al., 2015, 2016), and usable privacy (Chen et al., 2024). Specifically, we can empirically observe users’ interactions with Instagram Story ads, as well as inferring dwelling time over Story ads based on the first and last content change while a specific ad appears on screen. More generally, some broader questions we can answer with the dataset include: What empirical understanding of Instagram’s personalized advertisement algorithms can we obtain from quantitative data analysis? How do Instagram Story advertisements reveal the platform’s portraits of individual users’ preferences? And what are users’ folk theories of Instagram’s personalized ad algorithm?
Scenario 2: Uber Pricing
Ride-sharing platforms like Uber implement dynamic pricing strategies that adjust fare prices based on various factors including demand, time of day, and location. Collecting real-time price data for trip requests would allow researchers to study pricing algorithms, examine fare consistency across different user groups, and analyze how external factors influence ride-sharing costs. Previous research has attempted to understand these pricing mechanisms through qualitative measure (Lenjo, 2017) and quantitative analysis (Chao, 2019). However, current data sources cannot cover many user-specific aspects such as price variations across user groups and discount information. This data could help researchers investigate questions about algorithmic fairness in transportation access, analyze the relationship between surge pricing and local events, and study how pricing strategies affect user behavior in different geographical and temporal contexts. In our scenario, we focus on collecting the pricing information for participants’ Uber ride requests.
Scenario 3: Chrome Browser Discover Feed
Mobile browsers like Google Chrome integrate algorithmic news feeds into their search interface. These algorithmic feeds, often supported by deep learning, curates content based on users’ browsing behavior and inferred interests (Zheng et al., 2018). These algorithmic feeds, while work at an individual user level, have wide and profound social implications (Sandvig et al., 2014; Ribeiro et al., 2020).
Using Chrome as an example, accessing its Discover feed data from the user’s perspective enables research in algorithm auditing, helping to analyze content recommendation patterns, assess personalization mechanisms, and identify potential biases (Sandvig et al., 2014; Guess et al., 2023). Researchers have proposed audit studies to uncover discriminatory practices (Sandvig et al., 2014). Researchers can investigate how these algorithms shape information exposure by examining content diversity, frequency, and source representation. Understanding these dynamics in the Discover feed can identify biases impacting users’ information access and inform design improvements for more equitable content curation.
4.2. Study 1: Evaluating Crepe Usability with Researchers (In-Lab)
Study 1 evaluates how Crepe effectively supports researchers. While Crepe serves both researchers (collector creators) and users (data contributors), this study focuses specifically on the researcher experience during collector creation. We examine the usability of our Programming By Demonstration (PBD) approach in creating a data collector and try to understand what are researchers’ potential use cases of Crepe.
4.2.1. Procedure
Each one-hour session began with a brief introduction to Crepe’s motivation and core features. Participants then created data collectors for three representative usage scenarios (detailed in Section 4.1) using Crepe’s interface. After completing the tasks, participants completed a modified System Usability Scale (SUS) questionnaire (Brooke, 1996), including five-point Likert scale questions evaluating Crepe on its usability and utility for future research projects. We concluded with semi-structured interviews exploring participants’ envisioned uses of Crepe in their research. All studies were carried out in person and participants received a \$15 digital gift card for their time. To evaluate the accuracy of the PBD-generated Graph Queries, two authors independently analyzed whether each Query correctly captured the researcher’s specified data collection targets.
4.2.2. Participants
We recruited 5 HCI researchers through social media advertisements and word of mouth. All participants had experience with data collection and had interests in collecting users’ screen data for research purposes. Most participants have experience collecting various types of user data in the past (see participant demographics in Section A.2).
Figure 6. The usability questionnaire results for Study 1, evaluating researchers’ perception of Crepe’s usability and data collection capabilities. The tool received positive feedback, demonstrating its usability and utility for potential data collections studies.
[…]
A horizontal bar chart showing responses to 9 statements about Crepe’s usability on a 5-point Likert scale from ’Strongly agree’ to ’Strongly disagree’. Most responses trend positive, with ’Various functions are well integrated’ receiving the strongest agreement (5 respondents strongly agreeing). The chart uses a color scheme of dark blue for ’Strongly agree’ through pink for ’Strongly disagree’, with statement text aligned left and response counts displayed in bars.
4.2.3. Results
The usability evaluation showed positive results (Figure 6), with participants rating Crepe favorably on all questions. Most participants also showed a strong interest in using Crepe for future screen data collection studies and think it would not require much adaptation to Crepe.
In Study 1, as researchers created data collectors using Crepe, we also evaluated the programming by demonstration (PBD) pipeline for authoring candidate Graph Queries. Our evaluation was carried out on in total 15 created collectors (3 example scenarios for each of the 5 participants). We found that in all of the 15 occasions (15/15, 100%), our PBD pipeline successfully generated Graph Queries that correctly match users’ data collection goal (see Figure 4). Based on participant feedback, we iteratively improved the design of the user disambiguation process. We landed on our final design as described in Section 3.2.3, making it intuitive to pick the best Graph Query that matches researchers’ data collection goals.
Our participants identified several novel research opportunities enabled by Crepe’s screen data collection capabilities. P1 mentioned the interest in collecting Twitter (X)’s feed data to understand the platform’s feed algorithm. P2 would like to collect data from Uber to understand how ride prices change throughout a day and throughout different days of a week. P5 proposed to collect journal apps’ suggestion prompts to gauge what the app understands about users’ everyday activities. Interestingly, P1 also brought up the need to collect data based on certain conditions—for example, only collect social media feed under the “For you” tab instead of “Following”—which can further expand the utility of Crepe.
4.3. Study 2: Evaluating Graph Query Data Collection Accuracy (In-Lab)
Study 2 evaluates how accurately Crepe collects screen data in realistic app usage scenarios. While Study 1 focused on researchers’ experience creating data collectors, this study examines the technical performance of Graph Query in locating and extracting target data.
The nature of the study required us to conduct it as an in-lab study. Measuring the accuracy of Graph Query in data collection requires us to compare (1) the target data that showed up on users’ screens, and (2) the data Crepe is able to collect in the process. To obtain (1), we recorded participants’ screen during the in-lab study, to establish the baseline.
Metric Overall Instagram Uber Chrome
Recall 95.80% (506/528) 94.7% (269/284) 96.6% (115/119) 97.6% (122/125)
Precision 96.00% (504/525) 95.0% (268/282) 100% (115/115) 94.5% (121/128)
F1 Score 96.0% 94.8% 98.3% 96.0%
Table 2. The evaluation results for Crepe in the three app scenarios. Our in-lab study 2 demonstrates that Crepe performs well on on the three selected data collection scenarios. We discuss the observed edge cases that Crepe cannot handle well in Section 4.3.3.
[…]
This table presents performance metrics for a classification system across three digital platforms (Instagram, Uber, and Chrome) along with overall performance measures, where the total number of entries analyzed varies by platform with 528 entries overall, 284 entries for Instagram, 119 entries for Uber, and 125 entries for Chrome. The recall (true positive rate) measurements show 506 out of 528 (95.80%) overall, 269 out of 284 (94.7%) for Instagram, 115 out of 119 (96.6%) for Uber, and 122 out of 125 (97.6%) for Chrome. The precision (positive predictive value) results indicate 504 out of 525 (96.00%) overall, 268 out of 282 (95.0%) for Instagram, 115 out of 115 (100%) for Uber, and 121 out of 128 (94.5%) for Chrome. The F1 scores (harmonic mean of precision and recall) are 96.0% overall, 94.8% for Instagram, 98.3% for Uber, and 96.0% for Chrome, with Uber achieving perfect precision but slightly lower recall, while Chrome demonstrated the highest recall but the lowest precision, and Instagram maintained consistent performance across all metrics around 95%, resulting in strong overall performance across all platforms with F1 scores above 94%.
4.3.1. Procedure
We provided participants with a development phone with Crepe pre-installed and configured to collect data from apps in the three defined scenarios (Section 4.1). Each participant spent approximately one hour using these apps, allocating around 20 minutes per app. Participants were instructed to use each app naturally as they would in their daily life, while also feeling free to switch between apps when needed to mirror realistic usage patterns. All studies were carried out in person and participants received \$15 digital gift cards for each hour they participated.
To establish ground truth for evaluating collection accuracy, we simultaneously recorded the phone screen during the study sessions. An author labeled the screen recordings to identify all instances where target data appeared as the ground truth, and another author separately verified the results. We compared this ground truth data against the data collected by Crepe to compute collection accuracy.
4.3.2. Participants
We recruited 5 participants through social media advertisements and word-of-mouth. This provided approximately 240 minutes of naturalistic app usage data for our accuracy evaluation, generating a diverse set of UI states and interactions across different apps. Participant demographics and their app usage experience are detailed in Section A.3.
4.3.3. Results
Across all sessions, we recorded 528 data points across three apps as the benchmark (284 from Instagram, 119 from Uber, and 125 from Chrome). Crepe achieved an overall collection accuracy F1 score of 96.0% (precision: 96.0%, recall: 95.8%). Breaking down by app, collection accuracy F-1 scores were consistently high: Uber (98.3%), Instagram (94.8%), Chrome (96.0%). The minor variations in accuracy across apps can be attributed to differences in UI update frequencies and layout complexities.
Edge cases.
In the evaluation, we observed several scenarios where Crepe can make mistakes in data collection. First, some apps have UI layout variations for the same target data, requiring Crepe to set up multiple Graph Queries to collect the same data. In our evaluation, Instagram story’s ads are sometimes presented above a TextView containing “Sponsored”, and sometimes above a Button containing “Sponsored”. However, by demonstrating in both of these scenarios, researchers can still use Crepe to robustly handle such data collection tasks. In fact, we demonstrated three different UI layouts to Crepe during our evaluation study for Instagram Stories. Second, complex UI screens containing more than 500 UI elements can pose challenges to Crepe’s processing speed, potentially omitting data due to performance bottlenecks. We set a loose threshold to Crepe’s processing frequency to avoid repeatedly collecting the same data. We discuss some of our future improvement plans to address these challenges in Section 6.
In general, the results show that Graph Query can reliably collect screen data across various apps and UI screens. The collected dataset captured diverse scenarios including real-time price updates in Uber, dynamically positioned sponsored content in Instagram, and recommendation feeds in Chrome homepage. These results validate that Graph Query’s data collection mechanism is robust enough for real-world deployment while maintaining high accuracy across different apps and scenarios.
4.4. Study 3: Evaluating System Impact for Users (Field Study)
Now that we understand Crepe’s performance in lab, how well does it work in the wild? In Study 3, we conducted a field study with users who contributes their screen data through Crepe. Study 3 focuses on the real-world performance of Crepe, especially on its device performance and impact on data contributors’ regular phone usage experience. We designed Study 3 to evaluate how Crepe performs on different Android devices and OS versions, helping us catch any real-world deployment issues we couldn’t spot in the lab. Study 3 was conducted in example scenario 1 (Instagram story ads), since our target metrics will not significantly change when deployed for different apps’ data collections.
Participant Mobile OS Device Period Battery Usage
PC1 Android 14 Pixel 8 (own) 72 hours 1%
PC2 Android 14 Pixel 6 (provided) 72 hours 14%
PC3 Android 14 Pixel 6 (provided) 72 hours 9%
PC4 Android 13, OxygenOS 13.1 OnePlus 9R (own) 24 hours 0.91%
PC5 Android 12, Oxygen OS 12.1 OnePlus Nord (own) 24 hours N/A
PC6 Android 14 Pixel 6 (provided) 24 hours 2%
PC7 Android 9 Galaxy S10 (own) 24 hours 0%
Table 3. Mobile device usage data for study participants over 24 to 72 hour periods, including mobile OS, device model, and battery usage percentage. Three of the participants did not have an Android phone and used our provided device, with which they were asked to use Instagram following their regular habits during the study period. The participants decided the time period of the study based on their time schedule. Note that P5 reported that they could not found the battery usage item of Crepe on their phone Settings page after the collection. We were not able to help them find it either, despite our best efforts.
[…]
This table presents mobile device usage data for seven participants over study periods ranging from 24 to 72 hours. The table includes each participant’s mobile OS, device model, study duration, and the battery usage attributed to running the Crepe app. Three participants (PC2, PC3, and PC6) used provided devices (Pixel 6 with Android 14) because they did not own an Android phone. The remaining participants used their own devices, which varied in model and Android version. Battery usage varied by participant and device. PC1 used a Pixel 8 for 72 hours and reported 1% battery usage. PC2 and PC3, both using provided Pixel 6 devices for 72 hours, reported 14% and 9% usage, respectively. PC4 used a OnePlus 9R (Android 13, OxygenOS 13.1) for 24 hours and reported 0.91% battery usage. PC5, using a OnePlus Nord (Android 12, OxygenOS 12.1) for 24 hours, was unable to locate battery usage data in their device settings. PC6 used a provided Pixel 6 for 24 hours and reported 2% usage. PC7 used a Galaxy S10 (Android 9) for 24 hours and reported 0% battery usage. The results suggest that Crepe consumes minimal battery across a range of Android devices and OS versions. Most participants reported battery usage under 2%, indicating that Crepe is energy-efficient for background data collection. The only exception was PC2, who experienced 14% usage over 72 hours. A note below the table explains that PC5 could not find the battery usage stats despite assistance from the research team.